A warehouse employs 24 workers on first shift and 17 workers on second shift. Eight workers are chosen at random to be interviewed about the work environment. Find the probability that 3 first shift and 5 second shift workers are chosen.
Answer by stanbon(75887)

A warehouse employs 24 workers on first shift and 17 workers on second shift. Eight workers are chosen at random to be interviewed about the work environment. Find the probability that 3 first shift and 5 second shift workers are chosen.
--------
# of ways to succeed: 3*5 = 15
# of possible pairs: 24*17
-----
P(3 1st and 5 2nd) = 15/(24*17) = 5/(8*17) = 5/136
---------------------